# Batch Bayesian Optimization for High-Dimensional Experimental Design: Simulation and Visualization  [^1] A footnote for the title

### v1

@@ 3.1.2 Hartmann
font=footnotesize Refer to caption
Figure 5: BO results using UCB with $\beta$ = 1 on noise-free Hartmann function. Learning curve in (a) $||{\bf X}^{*}-{\bf X}_{\rm max}||$ and (b) $\mu_{\rm D}({\bf X}^{*})$ for 99 LHS BO models. The 25th percentile (green triangle), 50th percentile (red circle) and 75th percentile (blue triangle) regions are highlighted to exemplify poor, median, and good LHS BO models, respectively. (c) Visualization of the 3D representations for (x1 vs. x2) variables pair, for the 50th percentile BO model at the last iteration, showing how BO iterations zero in on ${\bf X}_{\rm max}$. Red circles are the 24 initial LHS selections and the blue to light green points are training points (blue progressively to green).(d) Parity plot of the 50th percentile BO model prediction vs. GT values for all 224 sampled points (LHS + 50 iteration at a batch size of 4).
The dependence of the BO results on the UCB exploration hyperparameter $\beta$ is shown in Table  S3. Similarly to Ackley, $\beta$ = 1 produces the lowest $\rm\langle IR(y)\rangle$ when using UCB policy. Comparing Table  S1 and  S3 for $\beta$ = 1 shows that ${\rm\langle IR}({\bf X})\rangle$ and $\rm\langle IR(y)\rangle$ are $\sim$ 210 times larger and $\sim$ 2 times smaller, respectively, for the Hartmann compared to the Ackley function. Our findings are in line with the benchmarking study of Liang et al.1
In Fig. 5(c), the ${\bf X}$ evaluated at each iteration of the 50th percentile (median outcome) LHS sampling are shown superimposed on the 3D representations of the GPR model. The red dots are LHS points and the blue-to-green dots represent BO learning progression. The projections onto the other pairs of input variables are shown Fig.  S6. The GPR model of the median sampling, hence all the percentile samplings above the median, accurately identify ${\bf X}_{\rm max}$ of the test function. The sampled points (blue to light green circles) converge on ${\bf X}_{\rm max}$ and no evaluated points appear near ${\bf X}_{\rm max,2}$ demonstrating that in this case, the BO model correctly identifies ${\bf X}_{\rm max}$. The parity plot in Fig. 5(d) compares the median BO model predictions with the GT data for all the evaluated points. RMSE value for this plot is $2.6\times 10^{-4})$. The normalized RMSE value is $7.8\times 10^{-5}$, which corresponds well to the converged square root value of the GNV hyperparameter (Supplementary Fig.  S7(h)).
Fig. S8 shows still snapshots of the BO iterations for the 50th percentile LHS sampling, projected onto the x1-x3 plane, whereas Movie (2) shows the entire iterative progression of the BO. Note that in Fig. 5(c) the 3D representation is of the GP model of the final iteration whereas in Fig. S8, the evolution of the GPR model is illustrated. Initially, the GPR model displays a very flat landscape without a peak and the $\mu_{\rm D}({\bf X}^{*})$ value is $<2$. As iterations progress, the GPR model develops a peak near the GT max ${\bf X}_{\rm max}$ (black cross). By the tenth iteration, the BO model correctly identifies the ${\bf X}_{\rm max}$. Note that there is no peak near the $2^{\rm nd}$ max ${\bf X}_{\rm max,2}$ (red star).
3.1.3 Comparing the UCB and the EI acquisition functions
The UCB and EI acquisition policies are compared for the Ackley and the Hartmann function optimization in Table 1.
@@ 3.1.3 Comparing the UCB and the EI acquisition functions
font=footnotesize Test Function Acquisition Policy $\rm\langle IR({\bf X})\rangle$ 1 ×$\times 10^{-2}$ $\rm\langle CR({\bf X})\rangle$ $\rm\langle IR(y)\rangle$ 1 ×$\times 10^{-2}$ $\rm\langle CR(y)\rangle$ Ackley UCB ($\beta$ = 1) 0.11 1.17 1.63 4.88 Ackley EI ($\xi$ = 0) 0.59 1.87 9.35 8.61 Hartmann UCB ($\beta$ = 1) 23.1 18.5 0.84 3.31 Hartmann EI ($\xi$ = 0) 33.0 18.9 0.91 3.10
Table 1: Comparison between UCB and EI acquisition policies with noise-free Ackley and Hartmann test functions. Exploration parameter settings ($\beta$ for UCB and $\xi$ for EI. respectively) are also listed.
It lists our four BO performance evaluation metrics that rely on instantaneous and cumulative regrets (${\rm\langle IR}({\bf X})\rangle$, ${\rm\langle CR}({\bf X})\rangle$, ${\rm\langle IR}(y)\rangle$, and ${\rm\langle CR}(y)\rangle$, the lower the value, the better the performance). The EI acquisition policy exploration hyperparameter $\xi=0$ was determined through optimization, similarly to the UCB acquisition function hyperparameter $\beta$. $\xi$ = 0 resulted in the lowest ${\rm\langle IR}({\bf X})\rangle$ on both test functions, as shown in Table  S2 and Table  S4. Independent of the metric or objective function, UCB yields lower regret values except for $\rm\langle CR(y)\rangle$ for Hartmann, where EI produces a slightly lower value (3.10) than UCB (3.31). The difference is observed in ${\rm\langle IR}({\bf X})\rangle$ for the Ackley function where the UCB value is more than 5 times lower. Figs. S9 and  S10 show the superior performance of UCB compared to EI visually via the evolution of $||{\bf X}^{*}-{\bf X}_{\rm max}||$ and $\mu_{\rm D}({\bf X}^{*})$, with the most significant difference for the Ackley test function. Overall, we conclude that UCB significantly outperforms the EI, especially in terms of the convergence of $\mu_{\rm D}({\bf X}^{*})$, which is also in line with Liang et al.1.
3.2 Noise
3.2.1 Comparing Utility functions in the presence of noise
Fig. 6a-d compares the learning curves using ${\rm Max}(y)$ (a, c) vs $\mu_{\rm D}({\bf X}^{*})$ (b, d) utility functions with 5% noise for both Ackley (a, b) and Hartmann (c, d) test functions. The slower convergence in Fig. 6 compared to Figs. 3 and 5 is due to the presence of noise in the data. The ${\rm Max}(y)$ results (left column) reveal that about 5 LHS samplings yield ${\rm Max}(y)>A({\bf X}_{\rm max})$ for Ackley, and the majority of LHS samplings produce ${\rm Max}(y)>H({\bf X}_{\rm max})$ for Hartmann. $y$ can exceed the GT maximum value because of the noise when evaluating ${\bf X}$ close to the maximum or second maximum. In such a case, the learning curve is tracking the outliers, i.e., data point with the largest noise value, and the problem is exacerbated at higher noise levels. In contrast, when $\mu_{\rm D}({\bf X}^{*})$ is used as the utility function in the BO, almost no LHS sampling result in $\mu_{\rm D}({\bf X}^{*})$ exceeding the GT maximum value. Furthermore, the $\mu_{\rm D}({\bf X}^{*})$ learning curve of the Ackley test function (Fig. 6b) reveals more clearly that the algorithm has not yet converged to the optimum compared to the ${\rm Max}(y)$ curve (Fig. 6a). Overall, we conclude that using $\mu_{\rm D}({\bf X}^{*})$ as the utility function for the learning curves is a more suitable measure of BO algorithm than Max(y), especially when noise is present.34 We have tracked $\mu_{\rm D}({\bf X}^{*})$ in all other learning curves in the paper.
[…]
font=footnotesize Refer to caption
3.2.2 Ackley
Fig. 7 (a), and (b), show the learning curves of the Ackley test function for noise levels of 2%, 5%, 7%, and 10%, from left to right, using EI as the acquisition policy. With 2% noise, the curves look similar compared to the no-noise case although at iteration 10, one LHS sampling does not satisfy $||{\bf X}^{*}-{\bf X}_{\rm max}||<10$ whereas all satisfied this criterion for the noiseless case. At 5% noise, we observe that many samplings do not reach the GT maximum but the 50th percentile sampling still reaches a value close to ${\bf X}_{\rm max}$. This indicates that a little more than half of the samplings find the global maximum location by the end of BO despite $\mu_{\rm D}({\bf X}^{*})$ is lower than $A({\bf X}_{\rm max})$.
[…]
font=footnotesize Refer to caption
Figure 7: Learning curve of $||{\bf X}^{*}-{\bf X}_{\rm max}||$ in top row (a) and $\mu_{\rm D}({\bf X}^{*})$ in the second row (b) for the Ackley function using LP as the batch-picking method and EI as the acquisition function for noise levels of 2% ($\xi$ = 0), 5% ($\xi$ = 0.05), 7% ($\xi$ = 0.1), and 10% ($\xi$ = 0.05) from left to right. For visualization, (c) 3D representations and (d) parity plots of the 50th percentile BO models for the same noise levels. In the 3D representations, x1 and x2 are projected with 224 learning points, including 24 initial LHS points (blue) and 50 iterations (from light to dark red) of BO points.
Exploration hyperparameters are investigated for both UCB and EI acquisition functions as a function of noise level, based on the performance metric ${\rm\langle IR}({\bf X})\rangle$ (Table  S9 and  S10). Comparing UCB $\beta=1$ and EI $\xi=0$, we find that for noise levels above 3%, the EI acquisition performs better than UCB. The better performance using the EI acquisition function motivated us to compare BO results for different noise levels using EI in Fig. 7. Neither EI nor UCB have been initially designed to be used for noisy objectives,44 and ${\rm\langle IR}({\bf X})\rangle$ are indeed high beyond 10% of noise, indicating most of the LHS samplings have not resulted in a converged BO. Lately, acquisition functions have been designed specifically for noisy objectives, such as Noisy-EI,44, 45 however, these have not yet been integrated into most BO packages widely used in materials science applications. Table  S11 shows the acquisition function and hyperparameter that yields the lowest ${\rm\langle IR}({\bf X})\rangle$ for each noise level, and reveals that for 2% noise, UCB with $\beta=2$ is optimal whereas for 5% and 10%, EI with $\xi$ = 0.05 is optimal, and for 7% noise, EI with $\xi=0$ is optimal. Generally, low $\xi$ values that lead to more exploitation perform clearly better in low-noise scenarios for Ackley, whereas in high-noise scenarios, the choice of $\xi$ does not affect the result drastically. In the case of UCB, the effects of the exploration hyperparameter on the convergence are not as drastic but the low exploration hyperparameter values also perform better in low-noise scenarios. Thus, based on this benchmark, low exploration choices are a robust option for EI and UCB with Ackley-type objectives.
Fig. 7(c) shows the GPR model of the sampling with the $50^{\rm th}$ percentile ${\rm\langle IR}({\bf X})\rangle$ and Fig. 7(d) shows the corresponding parity plots. For 2% noise, the GPR model resembles the GT; for 5%, the peak is slightly degraded; for 7%, the peak is significantly degraded; and finally, for 10%, the landscape is almost flat and the Ackley peak has all but disappeared. A zoom-in 2D heat map is shown in Fig. S15, and the GNV hyperparameters are shown in Fig. S16. Supplementary Movie (5) illustrates the progression of BO learning, showcasing the transformation from the initial to the final iteration of the $50^{\rm th}$ percentile sampling with 5% noise. Overall, we observe good performance at 2% noise with performance degraded as noise increases to almost no BO effectiveness at 10% noise. It is clear from 7(d) that the GPR model misses the peak when noise is above 5%. For 10% noise, the model was only covering $y$ values below -15, significantly lower than the peak. However, for Ackley, these regions constitute almost the entire ($\langle$ 99.99%) GT hypervolume. This fact is important when we compare BO performance for Ackley vs. Hartmann in Fig. 12.
3.2.3 Hartmann
[…]
font=footnotesize Refer to caption
Figure 8: Learning curve of $||{\bf X}^{*}-{\bf X}_{\rm max}||$ in top row (a) and $\mu_{\rm D}({\bf X}^{*})$ in the second row (b) of the Hartmann function using LP as the batch-picking method and EI as the acquisition function for noise levels of 2% ($\xi$ = 0.005), 5% ($\xi$ = 0.1), 7% ($\xi$ = 0), and 10% ($\xi$ = 0) from left to right. For visualization, (c) 3D representations and (d) parity plots of the 50th percentile BO models for the same noise levels. In the 3D representations, x1 and x2 are projected with 224 learning points, including 24 initial LHS points (red) and 50 iterations (from blue to green) of BO points.
We investigated the optimal exploration hyperparameters for both UCB and EI acquisition functions as a function of noise, based on the performance metric ${\rm\langle IR}({\bf X})\rangle$ (Table  S12 and  S13). Trends similar to the Ackley test function are found for Hartmann, albeit Hartmann test function is less sensitive to the choice. The smallest ${\rm\langle IR}({\bf X})\rangle$-yielding values are used for Fig.8 (EI with $\xi=0.005$, $0.1$, 00, 00 for 2%, 5%, 7%, and 10% noise, respectively). Table  S14 highlights the best-performing acquisition function and hyperparameter for Hartmann test function at each noise level. EI is the best acquisition function for five of the noise conditions whereas UCB is the best for six of the noise conditions with only small differences between each. This indicates that for the Hartmann landscape, UCB and EI have similar BO performance.
Fig. 9 and in Movie (6) (see supplementary) show the full iterative progression of the BO of the Hartmann function with 5% noise using the EI acquisition policy ($\xi$ = 0.1). Initially, the maximum of the GPR landscape is not located near ${\bf X}_{\rm max}$. However, at iteration 32, the optimum has been identified. In comparison, in the noise-free case (Fig.  S8), ${\bf X}_{\rm max}$ and $H({\bf X}_{\rm max})$ are identified after only 5 iterations.
[…]
font=footnotesize Refer to caption
Fig. 10(a) shows the fraction of LHS samplings where $||{\bf X}^{50}-{\bf X}_{\rm max}||<||{\bf X}^{50}-{\bf X}_{\rm max,2}||$. As noise increases, the fraction of LHS samplings that find ${\bf X}_{\rm max}$ decreases from 75% to 30%. The observation matches with Fig. 8, in which we studied noise up to 10% and observed that the 50th percentile still find ${\bf X}_{\rm max}$. However, further increase in noise results in more than 50% of the LHS samplings ending up closer to ${\bf X}_{\rm max,2}$ compared to ${\bf X}_{\rm max}$. This indicates that for a landscape with almost degenerate maxima, at high noise, the BO is not able to distinguish between the two. This could arise from the area of the convex peak of the second maximum being large compared to the first, as well as the relative locations of the two optima. Whenever the noise level is higher than the difference in the values of the true global optimum and other competing optima, BO has only little evidence for determining which one of the optima is better. However, the ability of BO to find an optimum diminishes more slowly with the increasing noise than the ability to find the global optimum.
[…]
font=footnotesize Refer to caption
Locations ${\bf X}^{50}$ are illustrated in Fig. 10(b-d) for 0%, 10%, and 20% noise levels. Without noise, all samplings result in ${\bf X}^{50}$ are clearly near either ${\bf X}_{\rm max}$ or ${\bf X}_{\rm max,2}$. But at 10% noise, a few samplings end up in between the two locations and are attributed to the first or second maximum only through the criterion $||{\bf X}^{50}-{\bf X}_{\rm max}||<||{\bf X}^{50}-{\bf X}_{\rm max,2}||$. The number of ”stray” samplings increases when noise increases to 20%.
3.2.4 Noise Overestimation Problem and Solution
[…]
font=footnotesize Refer to caption
@@ 3.2.4 Noise Overestimation Problem and Solution
Figure 11: Ackley learning curve of $||{\bf X}^{*}-{\bf X}_{\rm max}||$ (top row) and $\mu_{\rm D}({\bf X}^{*})$ (bottom row) for 10% noise: (a) and (b) of $\rm Max(y_{GT})$ and (c) and (d) of noiseless kernel amplitude. UCB is used as the acquisition policy.
@@
Therefore, we considered other options to incorporate noise in the benchmarking studies. We argue that the kernel amplitudes under noiseless conditions can be taken as the signal level. Thus, a more physical way to set the noise level is setting noise proportionally to the noiseless kernel amplitude rather than $\rm Max(y_{GT})$. This choice stemmed from the observation that the kernel amplitude plays a pivotal role in enhancing the precision of predictions for the objective function during the BO process.
@@
Fig. 11 compares 10% noise case incorporated in two noise frameworks for the Ackley function: proportional to $\rm Max(y_{GT})$ (a-b) and proportional to the noiseless kernel amplitude (c-d). The comparison for the Hartmann function is shown in Fig. S23. The noiseless kernel amplitude is equal to 0.192 and 0.184 for Ackley and Hartmann, respectively. For the kernel amplitude noise framework (Fig. 9(c-d)), all final BO models are well-optimized for the Ackley case. Conversely, for the $\rm Max(y_{GT})$ noise framework (Fig. 11(a-b)), only around 25% of the final BO models exhibit successful optimization, while the remaining fail to achieve satisfactory levels of optimization. The ${\rm\langle IR}({\bf X})\rangle$ value of the $\rm Max(y_{GT})$ is two orders of magnitude larger than that of kernel amplitude noise frameworks for the Ackley case: $31.7\times 10^{-2}$ and $0.38\times 10^{-2}$, reflecting the effect of noise over-estimation on BO outcome.
@@
3.2.5 Comparing Noise Effects on Ackley and Hartmann
@@
Fig. 12c shows that $\rm\langle RMSE\rangle$ values (averaged from 99 LHS BO models’ RMSE’s) for both functions increase with noise levels, with the two being approximately the same for $\langle$ 5% noise. At higher noise levels, $\rm\langle RMSE\rangle$ increases much more rapidly for Hartmann compared to Ackley. At first glance, this result is anti-intuitive since the GPR model at 10% noise level does not resemble the Ackley function at all (Fig. 7c), yet the RMSE value is lower than the Hartmann case where the GT landscape is clearly visible (Fig. 8c). This peculiar result is the direct consequence of the Ackley peak occupies only a miniscule fraction of the hypervolume. With the increasing noise, the GPR models completely missed the Ackley optimum region, but instead explore the plateau region where the GT values are small. The small $\rm\langle RMSE\rangle$ between the GT and GPR prediction indicates that BO correctly models the plateau region, and the large ${\rm\langle IR}(y)\rangle$ comes from missing the peak. Finally, the squre root of final model GNV’s averaged over all 99 LHS samplings, $\rm\langle\sqrt{GNV}\rangle$, is shown in Fig. 12d as function of noise. Up to 9% noise, $\rm\langle\sqrt{GNV}\rangle$ values are similar for both the Hartmann and Ackley functions, but beyond 9%, Ackley’s $\rm\langle\sqrt{GNV}\rangle$ becomes slightly lower. From 10% noise onwards, fitting the GPR GNV hyperparameter becomes more challenging for Ackley, while Hartmann’s $\rm\langle\sqrt{GNV}\rangle$ shows a stronger correlation with increasing noise levels, as expected.
@@ 4 Conclusion
We investigated high-dimensional Bayesian optimization (BO) and its relevant simulation and design choices in scenarios that involve noise. We have applied batch BO to the 6D Ackley function, representative of a needle-in-a-haystack problem, and the 6D Hartmann function, representative of a problem with a false maximum that are nearly degenerate with the global maximum.
The optimization progression is visualized with learning curves in both $\bf X$ and $y$, evolution curves of hyperparameters, 3D projections of the final Gaussian process regression (GPR) surrogate models and as a function of iterations, and parity plots. With noise, we show that using the optimal posterior mean of the model, $\mu_{\rm D}({\bf X}^{*})$, as the utility function is a more robust way to show BO convergence than using the value of the optimal objective, ${\rm Max}(y)$. These choices on how to track the BO convergence and ensure the proper progression of the optimization gain an increasing importance when BO tasks are performed in high-dimensional search spaces – where the surrogate model landscapes are difficult to visualize comprehensively.
In the absence of noise, BO is able to efficiently find the maximum of the Ackley GT function $({\bf X}_{\rm max})$ whereas for the Hartmann function, 30% of the LHS samplings ended in the false ground truth (GT) maximum $({\bf X}_{\rm max,2})$. We found that in the absence of noise, the UCB acquisition function with exploration hyperparameter $\beta=1$ yielded the best convergence in terms of the lowest instantaneous regret in $\bf X$, ${\rm\langle IR}({\bf X})\rangle$, compared to other values of $\beta$ or acquisition function EI with different values of the exploration parameter $\xi$.
We showed how BO performance is strongly degraded for the Ackley function by noise, with BO model not able to produce the GT maximum when 10% noise is present. For the Hartmann function, we also observe degradation but convergence to the optimum is still maintained up to 15% noise. With increasing noise, a larger fraction of the LHS sampling BO models ends at ${\bf X}_{\rm max,2}$. We inspected ${\rm\langle IR}({\bf X})\rangle$ normalized with respect to the range of the GT function as a function of noise. We found that Hartmann has a large ${\rm\langle IR}({\bf X})\rangle$ in the noiseless case because of the presence of the second maximum but the error increases relatively slowly with the increasing noise level, while ${\rm\langle IR}({\bf X})\rangle$ for Ackley is highly dependent on the noise level, increasing drastically above 4% noise level. These observations show that the specifics of the objective function, i.e., the optimization domain in the experimental case, affect the dependence of the BO performance on the noise, and should thus be evaluated in as much detail as possible when setting up experimental BO.
@@
We also show that simulating noise with respect to $\rm Max(y_{GT})$, as commonly done in the BO literature, is likely a significant over-estimation of the noise that would be seen in experiments, and that simulating noise with respect to the noiseless kernel amplitude of a GPR model fitted on the objective may represent more accurately signal-to-noise ratio in experiments. When using the noise relative to the noiseless kernel amplitude, BO is able to optimize the Ackley function at 10% noise in contrast to the $\rm Max(y_{GT})$ reference level. This highlights the importance of simulating noise in an realistic way prior to experiments, to be able to evaluate the required experimental budget and the feasibility of the optimization correctly.
Overall, we find the convergence of BO deteriorates with increasing noise level, but the behavior depends on the landscape of the optimization domains. Noise makes it harder to locate the global optimum in the Hartmann test function with a false maximum, but does not affect the objective value as much, whereas for the Ackley test function with a sharp peak, missing the location of the peak results in a complete loss in producing the function landscape. This work addresses the challenges of high-dimensional optimization, the importance of optimizing exploration hyperparameter under noisy environment, and introduces visualization techniques and performance metrics to track the progression of BO.

### v3

@@ 4.2.3 Landscape with Nearly Degenerate Maxima, Noisy Hartmann
4.2.4 Simulating Noise Amplitude to Better Resemble Experiments
Thus far in the benchmarking studies, we have incorporated noise as a percentage of ${\rm Max}(y_{\rm GT})$ for both test functions, as commonly done in the BO literature.[8, 52, 53, 50] Noise is thus incorporated as a Gaussian distribution with a zero mean and the standard deviation of the specified noise level (Eq. ((7))), which is added to the objective function value, $f({\bf X})$, according to Eq. ((8)).
While adding noise in this fashion is convenient in simulations, experimentally, noise is referenced to signals represented by the signal-to-noise ratio (SNR), which does not have a direct correspondence in the BO approach. First, when running simulated BO, noise incorporation proportionally to $\rm Max(y_{GT})$ can significantly overestimate the general noise level.[9] Secondly, since the benchmarks are utilized for making informed decisions for BO settings of future experiments, the noise level of the experiment to be used in the benchmarks is estimated based on repetitions of samples on a few points of the search space. In this case, the selected points are likely not to represent the true maximum since otherwise the experimental BO campaign would hardly be initiated in the first place. At the same time, noise proportional to $\rm Max(y_{GT})$ is relatively more detrimental for heterogeneous functions such as the Ackley than for smooth domains like the Hartmann, as seen in Figs. 7 and 8, which complicates the interpretation of the consequences. In experimental setups with multi-dimensional input parameters and complex objective functions, the over-estimation of the noise is generally considered safe. However, as a downside, the over-estimation of noise can lead to performing more experimental evaluations than necessary, increasing time and cost.
@@ 4.2.4 Simulating Noise Amplitude to Better Resemble Experiments
Figure 11: Ackley learning curve of ${\bf X}^{*}-{\bf X}_{\rm max}||$ (top row) and $\mu_{\rm D}({\bf X}^{*})$ (bottom row) for 10% noise: (a) and (b) of $\rm Max(y_{GT})$ and (c) and (d) of noiseless kernel amplitude. UCB is used as the acquisition policy.
@@
Therefore, we considered other options to incorporate noise in the benchmarking studies. We argue that the kernel amplitudes under noiseless conditions can be taken as the signal level. Thus, a more physical way to set the noise level is setting noise proportionally to the noiseless kernel amplitude, as expressed in Eq. (9), rather than $\rm Max(y_{GT})$ (Eq. (8)). This choice stemmed from the observation that the kernel amplitude plays a pivotal role in enhancing the precision of predictions for the objective function during the BO process.
@@
Fig. 11 compares 10% noise case incorporated in two noise frameworks for the Ackley function: proportional to $\rm Max(y_{GT})$ (a-b) and proportional to the noiseless kernel amplitude (c-d). The comparison for the Hartmann function is shown in Fig. S23. The noiseless kernel amplitude is equal to 0.192 and 0.184 for Ackley and Hartmann, respectively. For the kernel amplitude noise framework (Fig. 11(c-d)), all final BO models are well-optimized for the Ackley case. Conversely, for the $\rm Max(y_{GT})$ noise framework (Fig. 11(a-b)), only around 25% of the final BO models exhibit successful optimization, while the remaining fail to achieve satisfactory levels of optimization. The ${\rm\langle IR}({\bf X})\rangle$ value of the $\rm Max(y_{GT})$ is two orders of magnitude larger than that of kernel amplitude noise frameworks: $31.7\times 10^{-2}$ and $0.38\times 10^{-2}$, reflecting the effect of noise over-estimation on BO outcome.
@@
4.2.5 Effects of Noise Depend on the Problem Landscape
@@
Fig. 12c shows that $\rm\langle RMSE\rangle$ values (averaged from 99 LHS BO models’ RMSE’s) for both functions increase with noise levels, with the two being approximately the same for $\langle$ 5% noise. At higher noise levels, $\rm\langle RMSE\rangle$ increases much more rapidly for Hartmann compared to Ackley. At first glance, this result is anti-intuitive since the GPR model at 10% noise level does not resemble the Ackley function at all (Fig. 7c), yet the RMSE value is lower than the Hartmann case where the GT landscape is clearly visible (Fig. 8c).
This peculiar result is the direct consequence of the Ackley peak occupies only a miniscule fraction of the hypervolume. With the increasing noise, the GPR models completely missed the Ackley optimum region, but instead explore the plateau region where the GT values are small. The small $\rm\langle RMSE\rangle$ between the GT and GPR prediction indicates that BO correctly models the plateau region, and the large ${\rm\langle IR}(y)\rangle$ comes from missing the peak. Finally, the square root of the Gaussian noise variances in the final models averaged over all 99 LHS samplings, $\rm\langle\sqrt{GNV}\rangle$, is shown in Fig. 12d as function of noise. Up to 9% noise, $\rm\langle\sqrt{GNV}\rangle$ values are similar for both the Hartmann and Ackley functions, but beyond 9%, Ackley’s $\rm\langle\sqrt{GNV}\rangle$ becomes slightly lower. From 10% noise onwards, fitting the GPR GNV hyperparameter becomes more challenging for Ackley, while Hartmann’s $\rm\langle\sqrt{GNV}\rangle$ shows a stronger correlation with increasing noise levels, as expected.
@@ 5 Conclusion
Bayesian optimization (BO) is a machine learning method for effective optimization tasks; in experimental materials science, BO is useful for problems such as composition optimizations of mixed materials to achieve a desired but highly unusual property and process parameter optimizations to improve yield. In this work, we investigated high-dimensional BO — representing multi-variable optimization problems — and its relevant simulation and design choices in scenarios that involve non-negligible noise and performing work in batches, i.e., under conditions that are typical in materials experiments but not common in machine learning literature on BO.
We applied batch BO to the 6D Ackley function, representative of needle-in-a-haystack search problems such as searching for auxetic materials with a negative Poisson’s ratio,[13] materials with high thermoelectric figure of meric (ZT),[14] and high-entropy alloys that exhibit both high strength and high ductility,[15] and the 6D Hartmann function, representative of a problem with a false maximum that are nearly degenerate with the global maximum, exemplified by optimizing deposition parameters for perovskite solar cells,[16] searching for stable perovskite composition,[17], synthesizing silver nanoparticle by microfluidics,[18] and identifying print parameters for enhanced output quality.[19]
We developed a benchmark and visualization framework for planning and tracking high-dimensional BO, because problems above three dimensions are difficult to visualize and thus, controlling the high-dimensional optimizations is commonly challenged by difficulties to extract relevant information on its progress in a timely manner. We propose visualizing the optimization progress with (1) learning curves of both design variables $\bf X$ and objective variable $y$, (2) the evolution of hyperparameters, (3) 3D projections of the final Gaussian process regression (GPR) surrogate models as a function of BO iterations, and (4) parity plots representing the match of the experiment data and the surrogate model predictions.
We found that in the absence of noise, BO is not challenged by the needle-in-a-haystack type Ackley ground truth function and efficiently finds the input values associated with the global maximum $({\bf X}_{\rm max})$. In contrast, the Hartmann function sets a tougher challenge due to its local optimum $({\bf X}_{\rm max,2})$ close to ${\bf X}_{\rm max}$, with 30% of the LHS samplings ending up at ${\bf X}_{\rm max,2}$. This is reflected in the large instantaneous regret (${\rm\langle IR}({\bf X})\rangle$) value for Hartmann even without noise. In the absence of noise, the UCB acquisition function with exploration hyperparameter $\beta=1$ yields the best performance compared to other values of $\beta$ or acquisition function EI with any tested value of its exploration hyperparameter for both Ackley and Hartmann when evaluated according to ${\rm\langle IR}({\bf X})\rangle$, ${\rm\langle IR}(y)\rangle$ and cumulative regret ${\rm\langle CR}({\bf X})\rangle$. Only for ${\rm\langle CR}(y)\rangle$ in the Hartmann function does EI slightly edge out UCB.
We demonstrated that in noisy optimization scenarios as in typical materials science experiments, the progress of BO should be monitored via the predicted values of the objective function (the optimal posterior mean of the surrogate model, $\mu_{\rm D}({\bf X}^{*})$) rather than the best objective value collected (${\rm Max}(y)$). This is a more robust way to track the BO convergence because $\mu_{\rm D}({\bf X}^{*})$ is less susceptible to outliers. We showed that BO performance is strongly degraded for the Ackley function by noise, with BO not able to converge to the ground truth maximum when 10% noise is present. For the Hartmann function, we also observed degradation but convergence to the optimum is maintained up to 15% noise even though the convergence is increasingly distracted by the local optimum of the objective function ${\bf X}_{\rm max,2}$. This difference is reflected in our benchmark metrics: ${\rm\langle IR}({\bf X})\rangle$) for the Hartmann test function increases relatively slowly with the increasing noise level, while ${\rm\langle IR}({\bf X})\rangle$ for Ackley is highly dependent on the noise level, increasing drastically with noise above 4% level. These observations show that the specifics of problem landscape, i.e., the sensitivity of objectives on input variables, affect the BO performance in the presence of noise in experiments, and highlight the importance of domain expertise when adopting machine learning approaches for materials science research.
@@
As another factor relevant to planning experimental optimization campaigns in materials science, we showed that simulating noise with respect to the scale of the test function (here $\rm Max(y_{GT})$), as commonly done in the BO machine learning literature, might significantly over-estimate the noise in experiments. Fitting a GPR model on the objective and simulating noise with respect to the noiseless kernel amplitude of the same model likely represents more accurately the signal-to-noise ratio in experiments. In the latter case, BO is still able to optimize the needle-in-a-haystack Ackley function at 10% noise whereas in the former case it is not. This highlights the importance of simulating noise in a realistic way prior to experiments, to be able to evaluate the required experimental budget and the feasibility of the optimization campaign correctly.
This work addresses many important, but often-neglected, issues, such as which acquisition function to use or how to decide what value of exploration hyperparameter to use, in multi-variable optimization encountered in materials science research. In addition to testing BO under noisy experimental conditions, we introduce visualization tools and performance metrics to monitor optimization progress. This lays the ground work for moving BO from controlled studies toward robust, real-world adoption, bridging the gap between theory and experimental practice in materials research.
Funding
This work is supported by National Science Foundation CMMI-2135203. A.T. acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 01059891. J.W.P.H. and T.B. acknowledge the support from the Simons Foundation Pivot Fellowship. J.W.P.H. acknowledges the support of the Texas Instruments Distinguished Chair in Nanoelectronics.
Acknowledgements